### v1

# A Framework to Audit Email Address Privacy and Analyze Email Marketing Practices of Online Services and Apps

Abstract.
This study explores the widespread perception that personal data, such as email addresses, may be shared or sold without informed user consent, investigating whether these concerns are reflected in actual practices of popular online services and apps. Over the course of a year, we collected and analyzed the source, volume, frequency, and content of emails received by users after signing up for the 150 most popular online services and apps across various sectors. By examining patterns in email communications, we aim to identify consistent strategies used across industries, including potential signs of third-party data sharing. This analysis provides a critical evaluation of how email marketing tactics may intersect with data-sharing practices, with important implications for consumer privacy and regulatory oversight. Our study findings, conducted post-CCPA and GDPR, indicate that while no third-party spam email was detected, internal email marketing practices were pervasive, with companies frequently sending promotional and CRM emails despite opt-out preferences. The framework established in this work is designed to be scalable, allowing for continuous monitoring, and can be extended to include a more diverse set of apps and services for broader analysis, ultimately contributing to improved user perception of data privacy practices.
conference: ; ;
1. Introduction
Public perceptions of data privacy, particularly concerning personal data such as email addresses, can be significantly shaped by high-profile data breaches, feelings of powerlessness, and general misunderstandings. Isolated incidents of major data breaches contribute to the growing concern among users about the security of their personal information. For instance, the Yahoo data breach in 2013–2014 affected all 3 billion user accounts, exposing email addresses and other personal data (New York Times, 2017a). Similarly, the Equifax breach in 2017 compromised the personal information of approximately 147 million people, including email addresses and Social Security numbers (New York Times, 2017b). These incidents have heightened public awareness of vulnerabilities in corporate data handling practices. Such breaches exacerbate the concept of “digital resignation” where individuals feel powerless to control their personal information, accepting that their email addresses and other data may be shared or exposed without their consent (Draper and Turow, 2019). This resignation is further fueled by past corporate practices that shared data with third parties (Englehardt et al., 2018), before data privacy laws such as GDPR (Voigt and von dem Bussche, 2017) and CCPA (Pardau, 2018) were enacted. Hoofnagle et al. (Hoofnagle et al., 2012) found that users are often unaware of the extent to which their email addresses are collected and shared with third-parties, increasing feelings of vulnerability and mistrust. Furthermore, Acquisti et al. (Acquisti et al., 2015) found that even users who express privacy concerns can themselves engage in behaviors that compromise their email data, often due to a lack of viable alternatives or understanding.
A significant challenge for users is the difficulty of identifying which app or service may have leaked or shared their email address when they start receiving unsolicited spam or phishing emails. This issue arises because most users generally reuse the same email address across multiple online services and apps. Consequently, when unsolicited communications are received, users cannot determine the specific source responsible for the unauthorized sharing or leakage of their email information. This difficulty in identifying the source of leaked or shared email addresses is further exacerbated by pervasive email marketing practices, which were often connected to the sharing of email and other personal data with third parties. Companies often engaged in email marketing strategies that involve partnering with affiliates, utilizing data brokers, or participating in advertising networks to broaden their outreach (Sherman, 2021). These practices frequently entailed sharing or selling users’ email addresses, sometimes without informed consent or clear disclosure. As a result, users could receive unsolicited promotional emails from entities with whom they have no direct relationship, obscuring the boundaries of accountability. The interconnection between email marketing and email sharing practices not only increases the volume of unsolicited communications but also undermines users’ ability to trust that their personal information is handled responsibly.
After the introduction of data privacy laws like GDPR and CCPA, there still remains a significant gap towards systematically auditing online services and apps for their email address sharing and email marketing practices. Email marketing strategies have become increasingly sophisticated, often blurring the lines between legitimate marketing and unauthorized data sharing with third parties. Without robust auditing mechanisms, it’s challenging to assess compliance with these regulations or hold companies accountable for improper handling of email data. Developing such a framework is crucial for enhancing transparency and accountability in email marketing practices, ultimately improving user perceptions by rebuilding trust in how their personal information is managed.
Towards addressing this gap, our study offers a structured and scalable framework to assess how online services and apps handle user data. Our contributions include the following key steps and insights:
  • Audit of Online Services and Apps: We conducted a year-long audit of the top 150 popular online services and apps, opting out of any optional communication during the registration process. No third-party spam emails were detected within this time frame, suggesting that opting out during account creation is very effective in limiting unsolicited emails.
  • Email Marketing Analysis: Despite no unsolicited emails, we collected 4,847 emails received directly from these online services and apps. Our analysis revealed that companies still frequently used customer relationship management (CRM), alert, and promotional emails to engage users, even when opting out of promotional communications. This highlights the persistence of well-established marketing strategies, implemented internally.
  • Cross-Sector Insights: We identified significant differences in email marketing practices across sectors. Consumer industries, such as retail, heavily relied on promotional campaigns, while financial and service sectors prioritized CRM and alert emails. Some of these sector-specific behaviors offer insights into varying risks to data privacy.
These findings contribute to a comprehensive understanding of email marketing and data sharing practices, promoting transparency and accountability in how online services and apps handle user data.
2. Related Work
Email privacy has become a critical issue in the wake of increased data breaches and unauthorized sharing of personal information. Englehardt et al. (Englehardt et al., 2018) demonstrated the widespread use of email tracking, allowing third parties to gather user behavior data without consent. Their findings emphasize how email addresses, once shared, become tools for covert tracking, often leading to privacy violations. Similarly, Hoofnagle et al. (Hoofnagle et al., 2012) found that most users are unaware of how their email data is shared with third parties for advertising purposes, heightening concerns about privacy infringement. Plice et al.  (Plice et al., 2008) provided an economic analysis of unwanted commercial messages and their dependence on weak privacy controls. These studies highlight the necessity of open auditing mechanisms to protect email privacy.
Email marketing remains a central tool for customer engagement, but its effectiveness is tied to timing, frequency, and personalization. Thomas et al. (Thomas et al., 2022a) found that emails with lower levels of overt persuasion, such as CRM emails, tend to have higher open and click-through rates compared to promotional emails, aligning with the Persuasion Knowledge Model (Friestad and Wright, 1994). Similarly, Nuutinen (Nuutinen, 2015) emphasized the role of segmentation and personalization in improving email performance, noting that content relevance significantly impacts open and conversion rates. The importance of timing has also been explored extensively. Araujo et al. (Araújo et al., 2022) proposed a machine learning approach using regression models like Random Forest and LSTM to predict the optimal send times for marketing emails. Their study demonstrated that predictive algorithms could improve open rates by sending emails when users are most likely to engage, highlighting the importance of behavioral analytics in email marketing. Abrahams et al. (Abrahams et al., 2010), in their multi-industry study of large U.S. franchises, found that industries like hospitality and food service were prolific in email frequency, often sending 2-4 emails per week, compared to sectors like automotive, which sent emails infrequently. These findings provide valuable benchmarks for industry-specific marketing practices. Further research by Yang et al. (Yang et al., 2019) examined the impact of post-stay email marketing in the hotel industry, identifying how personalized, interactive emails increased customer retention and revisit intention. This aligns with the findings from (Budac, 2016), which highlighted the role of A/B testing, personalization, and frequency in determining email effectiveness across different industries.
Existing literature highlights the need for a balance between effective email marketing and safeguarding email privacy. While predictive models and personalized strategies can optimize engagement, they must be deployed ethically to avoid violating user privacy. Prior works emphasize that compliance with data protection laws, along with the adoption of transparent email marketing practices, is critical for maintaining consumer trust. Our study builds on this foundation by analyzing the email marketing and data-sharing practices of popular apps and services. Unlike prior works, which often did in-depth analysis of emails pertaining to a single company or industry type, our study instead focuses on a breath of services and apps across various sectors.
3. Experimental Setup and Data Collection
Our experimental setup is based on the assumption that if email addresses are leaked, shared, or sold by the services or apps we signed up for, those addresses would eventually start receiving spam or unsolicited emails from third parties. This assumption allows us to detect unauthorized sharing of user data through the presence of unsolicited communications, which are typically linked to marketing activities that involve purchasing user data from third-party vendors. It is important to note, however, that this approach has limitations. Specifically, we cannot detect instances where email addresses are leaked, shared or sold if those third parties do not subsequently send spam or other detectable communications.
To comprehensively evaluate the email sharing and marketing practices across different sectors, we selected a diverse set of widely used online services, mobile applications, and essential services, as follows:
  • Top 100 Online Services: We compiled a list of the most popular websites worldwide based on sources such as Cisco Umbrella [^1] https://umbrella-static.s3-us-west-1.amazonaws.com/index.html and Statista [^2] https://www.statista.com/. Services that were consistently ranked across the two sources were prioritized for inclusion in the final list.
  • Top 50 Mobile Applications: Similarly, a list of the top 50 mobile applications was compiled by analyzing app rankings from Google Play Store and Apple App Store. Notably, mobile apps that were already included in the top 100 online services (e.g., Facebook) were excluded to ensure non-redundant data. This ensures that our dataset captures a broader spectrum of both online services and apps across platforms.
Account Registration. After finalizing the list of online services and apps, we began the account registration process for each, utilizing standardized contact information. For each identified entity (online service and mobile app), a unique account was created following a standardized registration procedure. Unique email addresses were generated in the following format: [randomly-generated-name]$xxx$@[our-domain].[tld], with specific ranges of $xxx$ assigned as follows:
  • 000-099 for the 100 online services.
  • 100-149 for the 50 mobile apps.
This system facilitated clear differentiation between service types and ensured that each service and app had a unique identifier throughout the experiment. During registration, no additional subscriptions, email notifications, or promotional materials were selected. This was essential for maintaining consistency in data collection, allowing us to focus on analyzing unsolicited marketing emails from the service and app providers.
All received emails were collected in a single catch-all inbox, making it easier to manage, store, and scale the data collection. During the entire experiment, we did not click on any unsubscribe links in the received emails, which also helped maintain a consistent approach to data collection without influencing the outcome. At the end of the 361-day data collection period, all received emails were saved as EML files, containing the email headers, subject, body and any additional metadata, for our analyses.
Email Content Classification. To classify each email into one of the three classes (promotional, CRM, or alert), we employed Llama3.1, a leading, open source, large language model (LLM) given definitions directly from prior work (Thomas et al., 2022a). We utilized an in-context learning approach for this task, where the context includes the task description, the definitions from the prior work (the context), and the instance to be classified (Loukas et al., 2023). This approach has been shown to achieve human-level performance on sufficiently large LLMs with clear and concise rules (Brown et al., 2020). The full procedure is available in Appendix A. Understanding the distribution of these categories provides insights into how businesses strategically balance promotional campaigns, customer relationship management, and essential service notifications.
4. Findings
4.1. Descriptive Statistics
Over the 361-day collection period, we received 4,847 emails from 111 of the 150 services and apps, post registration (LABEL:table:all-domains in Appendix B). Of those emails received, 99.96% passed SPF checks and 81.64% passed DKIM checks. All emails that passed DKIM checks also passed SPF checks. Sender Policy Framework (SPF) verifies the sender’s IP address to ensure it is authorized by the domain owner, while DomainKeys Identified Mail (DKIM) adds a cryptographic signature to emails, verifying that the message hasn’t been altered in transit and is indeed from the claimed sender. Both mechanisms are crucial for protecting against spoofing and phishing attacks, which is why their high pass rates are significant in confirming the authenticity of the emails in our dataset. Post inspection of the emails received by their root domains and the email security checks via SPF and DKIM, we confirmed that the emails received were from the original companies’ domains with which we had made accounts and no third-party domain emails were detected in this dataset.
Initial descriptive statistics indicate that there is not a well-defined parametric distribution that determines the volume of emails received. The mean was 43 emails per domain with a median of 5, a skewness of 3.55 showing the heavy right tail, and kurtosis of 12.92. The data’s high skewness and kurtosis suggest a complex, highly skewed, and heavy-tailed nature that is not well captured by parametric distributions. Figure 1 shows that a few large senders make up the dominant amount of email volume received, with the cumulative distribution following the 80-20 Pareto principle, where the total volume of emails received are dominated by a few senders. Just the top 10 root domains make up 63.23% of the total volume.
Figure 1. Pareto diagram of total email volume.
To classify the content of the emails, we rely on definitions provided by (Thomas et al., 2022a), where emails were classified as either “promotional”, “CRM” (customer relationship management), or “alerts”. These email types convey distinct messages with differing levels of persuasion. It is the appropriate combination of these different types of messages at the right time that drive click-through and eventual conversion, or conversely, abandonment, by the potential customer. To classify the emails that we collected, we used an in-context classification approach with a general purpose LLM (Llama3.1-8b-instruct (AI@Meta, 2024)) with structured output. The prompt and structured output schema is provided in Appendix A. There were 3,346 (69.0%) promotional emails, 1,393 (28.7%) CRM emails, 92 (1.89%) alert emails, and 16 (0.33%) that were unable to be classified as they were unparseable.
The emails received do not just vary in the volume sent by companies, but also in the content of the messages being sent. Previous literature shows that strategic email marketing campaigns differ in both what the sender says (the content) and when they say it (temporal dimension) (Thomas et al., 2022b). Additionally, previous work has shown that not all companies engage equally in strategic email marketing (Abrahams et al., 2010), with some being more sophisticated actors than others. Therefore, the following subsections of the analysis are structured as such: (1) We explore a unique contribution of our dataset to the literature on this topic, which is the sender’s IP address and associated information about that sender through its Autonomous System Number (ASN). (2) A temporal analysis to show at what times and with what frequency emails are sent. (3) Clustering of companies based on their emails content and temporal patterns to determine where companies fall on their level of sophistication or general tactics.
Figure 2. Sankey diagram of the top 50 flows.
4.2. Senders by Autonomous System
We captured the IP address of the sender of each email, with that we then ran a lookup of the registered autonomous system for each of those IP addresses and were able to tie it back to the company that sent it. Because of this, we are able to identify that most of the emails received come from a third-party intermediary that specializes in business-to-business marketing and CRM services like Salesforce and Mailgun. Like the overall email volume, the email volume by ASN also follows the Pareto principle, where a few large providers are responsible for the sending of the large majority of emails. Specifically, the top twenty percent, 8 ASNs, produced 89.35% of the total email volume. Figure 2 shows the evidence that most of the email marketing flow originates from a few large, third-party providers.
However, understanding which companies use which email marketing service, or whether they use a third-party service at all, is not straightforward. Some companies send emails originating from their own registered ASN such as wish.com (Wish is the “doing-business-as” name for ContextLogic), Alibaba Group, LinkedIn, and Adobe (Adobe acquired Omniture as part of their web analytics platform). Others use specialized third parties, chiefly Salesforce, Mailgun, Sendgrid, and Sparkpost. A third group had emails originate from cloud service providers like AWS, GCP, and Rackspace. For those that originate from cloud service providers, using only the IP addresses we cannot ascertain what entity is managing the email campaigns at the application layer.
4.3. Temporal Analysis
The temporal dimension of this analysis examines both the timing (time of day and day of week) and the frequency of emails sent. First, we treat the incoming messages as a time series, decomposing it to understand the general trend and identify hypothesized seasonality as shown in Figure 3. Since account creation dates differ across services and apps, we standardize the series by the number of days since the first email received, which coincides with the account creation date for each app. This standardization provides a uniform time scale for the analysis.
Applying a Fourier transform to the combined time series revealed dominant frequencies, identified at the $2\sigma$ level, corresponding to various periodic cycles: annual, semi-annual, quarterly, and weekly. The identified frequencies ($f$), magnitudes ($|X(f)|$), and periods ($T$) are as follows:
  • Annual cycle: $f=0.0028$, $|X(f)|=834.38$, $T\approx 362$ days
  • Semi-annual cycle: $f=0.0055$, $|X(f)|=479.22$, $T\approx 181$ days
  • Quarterly cycle: $f=0.0138$, $|X(f)|=308.50$, $T\approx 72.4$ days
  • Weekly cycle: $f=0.1436$, $|X(f)|=393.39$, $T\approx 6.96$ days
The most relevant frequency for our analysis is the weekly cycle ($\approx 7$ days), denoted by $f=0.1436$ with a magnitude of $|X(f)|=393.39$. This weekly periodicity corresponds closely with the expected operational and behavioral cycles observed in email-sending patterns, reflecting regular weekly activity. We use the 7-day period as the seasonality figure for the time series decomposition. The decomposition highlights a linear, decreasing trend in the number of emails received over time. Additionally, it shows that the weekly seasonal component explains 1.17% of the variance in the number of emails received. The amount of seasonality in email sending behavior is mediated at the company level and will be explored in greater detail in following sections on the cross-sector and company-level analysis.
Figure 3. Additive time series decomposition of aggregate emails received.
Second, given that the observed weekly seasonality in sending patterns we look to specify the general time of day and day of week that email messages are sent. The polar heat maps shown in Figure 4 show that the most popular times messages are sent occur during the mid-week, Tuesday through Thursday, and emails are predominantly sent around midnight or during the afternoon work hours between 1300 and 1500 hours.
Figure 4. Email frequency by hour of day for each day of the week.
4.4. Clustering
The next task is to determine which companies exhibit similar email sending patterns, and then map those patterns to levels of sophistication from prior literature. To accomplish this unsupervised learning task, we first perform a principal component analysis (PCA) on all the features created from the previous analysis. Then, we perform K-Means to assign each company to a cluster.
The analysis began by loading the relevant email and ASN data and selecting the temporal and behavioral features.
  • Hourly Email Frequency: We computed the count of emails sent by each company per hour of the day (0-23), aggregated across all days.
  • Weekly Seasonality: We analyzed the weekly email sending pattern by determining the frequency of emails sent on each day of the week (0 = Monday, 6 = Sunday).
  • Professional Marketing Indicator: Each ASN description was checked against known marketing services (e.g., Salesforce, Mailgun, SendGrid, SparkPost) to create a binary feature indicating the presence of professional marketing practices.
  • Email Content Classification: We captured the distribution of email content (e.g., promotional, CRM, alert) for each company, allowing us to profile their communication style.
  • Total Email Volume: The overall number of emails sent by each company was computed to gauge their communication scale.
These features were combined into a comprehensive feature set, standardized using StandardScaler, and subsequently reduced to five principal components using PCA. The dimensionality reduction to five principal components retained over 80% of the variance in the data. The PCA results show the factor loadings and below is an interpretation of the PCA component loadings:
  • Component 1 captures overall email activity and promotional focus. Companies with high scores on this component send a high volume of emails, particularly promotional content, consistently across weekdays. This suggests a strategy that emphasizes frequent promotional outreach.
  • Component 2 represents email sending times, particularly in the morning hours. High scores on this component indicate companies that strategically send emails during morning hours (7 AM to 10 AM), possibly to increase open rates and customer engagement.
  • Component 3 reflects customer engagement through CRM and alert emails sent throughout the day and evening. Companies scoring high on this component focus on personalized communication and customer relationship management, engaging customers with CRM and alert messages at various times.
  • Component 4 emphasizes afternoon and evening email sending patterns. High scores on this component suggest companies that target specific times of the day, such as afternoon and evening hours, to optimize their communication strategies.
  • Component 5 highlights alert emails sent during non-standard hours, such as early morning and late evening. Companies with high scores on this component may send time-sensitive communications, such as urgent updates or notifications.
K-Means clustering was employed to group companies based on their PCA-transformed features. To determine the optimal number of clusters, we evaluated cluster validity using the Silhouette Score, a metric that measures how similar a point is to its assigned cluster compared to other clusters. We tested a range of cluster numbers (2-10) and identified an clustering configuration based on Silhouette Score and prior domain knowledge. For this analysis, a final model with five clusters was selected, as it effectively balanced intra-cluster coherence with inter-cluster separation. Using the PCA loadings, we interpret the characteristics of each cluster to better understand the distinct email marketing behaviors exhibited by companies within each group. The average scores of each cluster on the principal components highlight the dominant features that distinguish one cluster from another.
  • Cluster 0: Low-Activity Generalists
    Cluster 0 exhibits negative scores on Component 1 and slight negative or near-zero scores on other components. This cluster represents companies with generally low email activity, characterized by a lack of pronounced patterns in sending frequency, content, or professional marketing service usage. The absence of significant influence from any principal component suggests that these companies do not heavily invest in sophisticated or high-volume email marketing strategies. Their approach may be minimalistic or highly targeted with low engagement, reflecting a more traditional or less dynamic email marketing approach.
  • Cluster 1: Ultra High-Volume Promotional Marketers
    Cluster 1 stands out with a highly positive score on Component 1 and a negative score on Component 2. This combination indicates companies that engage in extremely high-volume, promotional-focused email marketing but may not prioritize morning sending times. The negative loading on Component 2 suggests these companies might not strategically time their emails for the morning hours. These firms dominate in terms of sheer volume, indicating a strategy that prioritizes broad promotional outreach, likely at the expense of targeted or personalized engagement.
  • Cluster 2: High-Volume Morning Promotional Marketers
    Cluster 2 is characterized by strong positive scores on Components 1 and 2 but negative scores on Components 3 and 4. This profile points toward companies that not only engage in frequent email activity but also strategically send emails during morning hours. The negative scores on Components 3 and 4 suggest less emphasis on CRM communications and afternoon or evening sends. These companies may leverage professional marketing services extensively to optimize their outreach.
  • Cluster 3: Diversified Marketers with Emphasis on CRM and Alerts
    Cluster 3 shows positive scores across all principal components, with particularly high scores on Components 2 and 3. These companies are distinguished by a strong focus on diversified email strategies, including CRM and alert-type communications sent throughout the day. The positive scores on Components 2 and 3 suggest these companies engage in targeted messaging at various times, aligning with customer-specific triggers or personalized engagement strategies. This cluster represents sophisticated marketers who prioritize customer engagement alongside promotional activities.
  • Cluster 4: Balanced Marketers with Strategic Timing
    Cluster 4 presents a moderate positive score on Component 1 and slight positive scores on Components 3, 4, and 5. This cluster captures companies that maintain a balanced approach to email marketing, with strategic timing and a mix of promotional and CRM-focused content. They are characterized by neither extremely high volume nor extreme specialization in any one type of content, suggesting a versatile and adaptive strategy. This cluster represents firms that aim for a well-rounded communication style, balancing frequency, content, and timing to optimize customer engagement without overwhelming recipients.
Figure 5. Cluster membership by top two principal components.
The clustering analysis, visually depicted in Figure 5, reveals distinct segments of companies based on their email marketing strategies, ranging from high-volume promotional tactics to CRM-focused approaches. The PCA loadings provide a clear view of the factors driving these segmentations.
  • Volume vs. Personalization Trade-off: Clusters 1 and 3 illustrate the trade-off between volume and personalization, with Cluster 1 companies prioritizing frequent promotional outreach, while Cluster 3 focuses on targeted CRM communication and customer engagement.
  • Professional Marketing Services Usage: Clusters 2 and 4 emphasize the strategic use of professional marketing tools, leveraging advanced platforms to optimize timing and content, indicative of a more sophisticated approach to email marketing.
  • Strategic Timing: Clusters 2, 3, and 4 highlight companies that not only consider what to send but also when to send, using timing as a strategic lever in their marketing mix.
4.5. Cross-Sector Differences
To explore potential differences in email sending behaviors across various sectors, we analyzed key metrics such as the total volume of emails sent, the types of emails (promotional, CRM, alert), and the timing of email sending (hourly and daily patterns). The Health sector was excluded from the analysis due to having only one company. Descriptive statistics for each sector were calculated to understand the average email sending behaviors (Table 1).
Table 1. Email volume and type by sector (averages).
Sector Total Promo CRM Alert
Consumer Discretionary 58.14 48.52 9.06 0.37
Financials 53.50 0.50 40.00 13.00
Communication Services 19.63 3.37 14.84 1.26
Information Technology 17.88 8.96 8.54 0.33
Email Volume and Type by Sector. A chi-squared test of independence was conducted to examine the relationship between email type and sector (Table 2). The test revealed a statistically significant association between email type and sector ($\chi^{2}(6)=1411.29,p<0.0001$), indicating that the distribution of email types varies by sector.
Table 2. Email types per sector.
Sector CRM Alert Promotional
Consumer Discretionary 571 23 3057
Communication Services 282 24 64
Information Technology 205 8 215
Financials 80 26 1
The significant chi-squared test suggests that sectors differ in the types of emails they send:
  • The Consumer Discretionary sector, which includes companies like wish.com, wayfair.com, and kohls.com, predominantly sends promotional emails (83.73% of their emails). This aligns with strategies to stimulate immediate consumer purchases through promotions.
  • The Financials sector focuses heavily on CRM (74.77%) and alert emails (24.30%), reflecting the importance of customer relationship management and timely notifications in financial services.
  • The Communication Services sector sends a higher proportion of CRM emails (76.22%), suggesting an emphasis on maintaining customer relationships and providing service updates.
  • The Information Technology sector exhibits a more balanced distribution between promotional (50.23%) and CRMemails (47.90%), indicating a dual focus on marketing products and supporting existing customers.
Variability Among Individual Companies. A one-wayANOVA was conducted to compare the average emails sent per company by sector. The results indicated no statistically significant difference ($F(3,103)=1.4487,p=0.2330$), suggesting that the average email volume per company is similar across sectors. Similarly, the Kruskal-Wallis H tests for both hourly and daily email sending patterns revealed no statistically significant differences between sectors. This indicates that sectors do not differ significantly in terms of when they send emails.
The lack of significant differences in email volume and timing between sectors implies that we cannot generalize about an individual company’s email marketing sophistication or patterns based solely on its sector. Individual companies within the same sector may adopt diverse strategies in terms of email frequency and timing, reflecting company-specific marketing approaches.
However, the significant differences in email types suggest that sectors have characteristic focuses in their email content, even if the overall volume and timing are similar. This means that while the sector provides some insight into the nature of email communications (e.g., promotional vs. CRM focus), it does not predict how intensively or when individual companies will engage in email marketing.
Moreover, although we did not find broad evidence of sector-wide differences in email volume or timing, it is noteworthy that the most prolific email senders are from the Consumer Discretionary sector, particularly retail and product-oriented businesses. Companies like Wish, Wayfair, and Kohls are among the highest volume emailers, emphasizing promotional content to drive sales.
The statistical analysis indicates a significant association between email type and sector, highlighting that different industries utilize email communications to achieve their specific business objectives:
  • Retailers in the Consumer Discretionary sector engage in high-volume promotional emailing, aiming to drive sales through frequent promotions.
  • The Financials sector prioritizes CRM and alert emails, essential for customer engagement and timely financial updates.
  • The Communication Services sector focuses on CRM communications, possibly to maintain service satisfaction and customer loyalty.
  • The Information Technology sector maintains a balanced approach, combining promotional efforts with customer support communications.
These insights suggest that while the overall volume and timing of emails are consistent across sectors, the content and purpose of emails are tailored to industry-specific strategies.
5. Discussion
This study provides a comprehensive analysis of email marketing strategies employed by the most popular online websites and mobile applications. By collecting emails where each unique email address corresponds to a single company, we have ensured a precise lineage and isolation of source company to end user. This methodological approach, combined with a year-long email data collection period encompassing 111 unique companies, offers both breadth and depth to our dataset, distinguishing our work from prior studies.
Our research contributes to the ongoing conversation about how companies utilize digital channels to market to and maintain relationships with their customers. The extensive dataset allows us to observe real-world practices over an extended period, providing insights into not only the frequency and content of communications but also the strategies underlying them.
The findings reveal no evidence of spam emails from third-party domains and a significant portion of consumer concerns regarding excessive promotional emails can be mitigated by opting out of marketing communications where possible. The data also shows that the bulk of unsolicited first-party emails originate from a small subset of companies, supporting the Pareto principle where approximately 80% of the emails come from 20% of the senders. This concentration suggests that aggressive email marketing is not a universal practice but is driven by specific companies, often within certain sectors like retail and consumer services.
ASN Mapping Shows Reliance on Third Parties. A novel aspect of our study is the analysis of email provenance via the originating Autonomous System Numbers (ASNs). By tracing the sender IP addresses to their corresponding ASNs, we uncover the reliance of companies on third-party marketing and Customer Relationship Management (CRM) services. Major intermediaries such as Salesforce, Mailgun, SendGrid, and SparkPost are identified as key players facilitating the bulk of email marketing traffic.
This approach sheds light on the ecosystem of email marketing infrastructure, highlighting how companies leverage specialized external platforms to execute their communication strategies. A complete picture, however, is elusive as many of the emails originate from public cloud providers such as AWS and GCP, so what third party is operating its service hosted on these hyperscalers isn’t able to easily be determined just from the IP address. It also raises questions about data privacy and the extent to which customer information is shared with third-party providers, a topic that warrants further investigation.
The “What” and “When” Matter. Our temporal analysis indicates that while the overall volume and timing of emails are consistent across sectors, the content and purpose are tailored to industry-specific strategies. For instance, the Consumer Discretionary sector predominantly sends promotional emails, aligning with their goal of stimulating immediate purchases. In contrast, sectors like Financials focus on CRM and alert emails, reflecting the importance of customer engagement and timely information dissemination.
The identification of distinct clusters through principal component analysis (PCA) and K-means clustering reveals that companies adopt varied email marketing tactics ranging from high-volume promotional outreach to personalized CRM communications. This diversity underscores the need for nuanced strategies that consider both the timing and content of emails to effectively engage customers.
5.1. Legal and Ethical Considerations
The findings raise important legal and ethical considerations regarding compliance with privacy policies and marketing preferences. Despite regulations like the CAN-SPAM Act in the United States and the General Data Protection Regulation (GDPR) in the European Union, consumers continue to report receiving unsolicited emails. Our study suggests that opting out during registration can reduce the volume of such emails, but also highlights that some companies may not fully adhere to stated preferences.
Aggressive marketing practices, especially those that aim to “wear down” consumers through persistent communications, can erode trust and infringe upon consumer rights. Prior research has demonstrated that excessive emailing can lead to customer fatigue and negative brand perception (Cao and Sun, 2019). Companies must balance their marketing objectives with respect for consumer autonomy and privacy.
Policymakers and regulatory bodies may find these insights valuable in assessing the effectiveness of current regulations and identifying areas where enforcement or additional guidelines are necessary. Our study identifies that a vast majority of marketing emails are facilitated by third-party services such as Salesforce, Mailgun, and SendGrid, with 89.35% of emails being sent through the top 8 Autonomous System Numbers (ASNs). This centralization indicates potential focal points for regulatory oversight. Additionally, despite existing regulations like the CAN-SPAM Act and GDPR, our findings suggest that some companies may not fully comply with stated privacy policies and marketing preferences, as evidenced by the persistent high volume of unsolicited promotional emails from certain sectors. Policymakers should consider these patterns when evaluating regulatory frameworks, ensuring that opt-out mechanisms are straightforward, transparent, and honored in practice. Enhancing regulations to specifically reference and include the use of third-party marketing services and enforcing compliance can further improve consumer experiences and uphold consumer rights.
5.2. Limitations and Future Work
The richness of our dataset extends beyond metadata to include the full subject lines and body content of emails. This presents opportunities for future research into the specific language, persuasive techniques, and thematic elements employed in email marketing. To do this, future work can explore the content of emails in greater detail through natural language understanding (NLU) techniques. Additionally, most of the emails in the dataset contain images that could be rendered and then analyzed, yielding insights into specific persuasion techniques that are employed. Moreover, our dataset can support studies on the effectiveness of different email strategies by correlating communication patterns with customer engagement metrics, where available. Finally, investigating the role of cultural, regional, or demographic factors in email marketing practices also presents an avenue for further study. Future work could also investigate whether smaller companies, which may face less regulatory scrutiny compared to popular online services and apps, are more likely to share or leak users’ email addresses without consent.
A key component of our framework is its low cost of ongoing maintenance and scalability. Once deployed, it can monitor any number of online services and apps passively. A key advantage of this is that we can detect any future changes in the inbound email messages. So for example, if a company in the monitored set has a breach that leaks customer email addresses and then those addresses are used for spam campaigns or sold to unauthorized third parties, our system would be one of the first authoritative detection sources of such a breach.
6. Conclusion
This study offers a detailed examination of email address privacy and marketing practices among leading online services and apps, highlighting key patterns in communication strategies and the use of third-party services. Our findings contribute to a better understanding of how companies balance marketing objectives with consumer preferences and the implications for trust and engagement in the digital marketplace. By shedding light on these practices, we aim to inform both industry and policy discussions, promoting approaches that respect consumer rights while enabling effective marketing communications.
Appendix A LLM Classification Workflow
The following code demonstrates the LLM classification workflow used to categorize emails as promotional, CRM, or alert.
1tagging_prompt = ChatPromptTemplate.from_template(
2 """
3You are an email classifier system. Your task is to classify the following email message as either promotional, CRM, or alert. Respond only with the JSON formatted classification based on the provided schema below, containing only the classification results and not the schema itself. Do NOT output the schema or any definitions-only provide the classification response.
4
5Here is the JSON schema of the response you must strictly follow:
6
7{schema}
8
9Here are the classification definitions (these are for your reference only, do not include them in the output):
10
11promotional: Emails offering price discounts intended to induce short-term actions such as purchasing. These are overt selling attempts with a call to action.
12
13CRM: Emails that engage, inform, or enhance the company’s relationship with the customer. These emails are not overtly sales-focused but aim for longer-term goals, like brand building or engagement.
14
15alert: Emails containing various notifications, updates, or alerts based on consumer preferences. These are informational and do not directly attempt to sell.
16
17Classify the following email strictly based on these definitions:
18
19Email:
20{input}
21"""
22)
23
24# Create the structured output format class
25class Classification(BaseModel):
26 sentiment: str = Field(..., enum=["promotional", "CRM", "alert"])
27 confidence: int = Field(
28 ...,
29 description="describes how confident the classification is, the higher the number the more confident in the classification",
30 enum=[1, 2, 3, 4, 5],
31 )
32 rationale: str = Field(description="The primary driver of the classification based on the definitions")
Listing 1: LLM classifier prompt and code.
Appendix B List of Domains
LABEL:table:all-domains shows the details and distribution of emails received from 111 different root domains.
Table 3. Distribution of emails received from 111 different root domains.
Root Domain Sector Cluster Total Emails Promotional CRM Alert
adobe.com Information Technology 0 37 24 13 0
airtable.com Information Technology 0 2 0 2 0
aliexpress.com Consumer Discretionary 0 37 33 3 1
apartmentfinder.com Consumer Discretionary 0 1 0 1 0
apartments.com Consumer Discretionary 0 8 1 7 0
autozone.com Consumer Discretionary 0 13 7 5 1
bandcamp.com Consumer Discretionary 0 4 1 3 0
bestbuy.com Consumer Discretionary 2 391 376 12 3
bloomberg.com Financials 0 8 0 5 3
bloombergbusiness.com Financials 4 99 1 75 23
booking.com Consumer Discretionary 0 2 0 2 0
canva.com Information Technology 0 3 1 2 0
capcut.com Communication Services 0 69 19 49 1
cars.com Consumer Discretionary 0 15 6 9 0
cnet.com Information Technology 0 1 0 1 0
codepen.io Information Technology 0 49 1 47 1
codewars.com Information Technology 0 3 0 3 0
coursehero.com Consumer Discretionary 0 22 7 15 0
discogs.com Communication Services 0 8 0 7 1
discord.com Communication Services 0 3 0 3 0
doordash.com Consumer Discretionary 0 3 3 0 0
dunkinrewards.com Consumer Discretionary 0 70 61 8 1
ebay.com Consumer Discretionary 0 70 57 13 0
etsy.com Consumer Discretionary 4 196 115 81 0
expedia.com Consumer Discretionary 0 28 24 3 1
flickr.com Communication Services 0 44 13 31 0
foodnetwork.com Consumer Discretionary 0 1 0 1 0
foxsports.com Consumer Discretionary 4 111 39 63 9
fragrancenet.com Consumer Discretionary 0 50 41 9 0
freecodecamp.org Information Technology 0 1 0 1 0
geeksforgeeks.org Information Technology 0 12 7 5 0
github.com Information Technology 0 2 0 1 1
google.com Communication Services 0 3 0 2 1
hackerrankmail.com Information Technology 0 4 0 4 0
homedepot.com Consumer Discretionary 0 1 0 1 0
hotels.com Consumer Discretionary 0 89 72 17 0
imdb.com Consumer Discretionary 0 2 0 2 0
instagram.com Communication Services 0 14 0 13 1
kayak.com Consumer Discretionary 0 3 0 3 0
kbb.com Consumer Discretionary 0 1 0 1 0
khanacademy.org Information Technology 0 49 21 28 0
kohls.com Consumer Discretionary 3 335 330 5 0
leetcode.com Information Technology 0 8 5 3 0
linkedin.com Communication Services 0 31 1 30 0
lowes.com Consumer Discretionary 4 411 404 6 1
lyftmail.com Consumer Discretionary 0 7 1 6 0
microsoft.com Information Technology 0 3 0 3 0
nationalgeographic.com Consumer Discretionary 0 1 0 1 0
newegg.com Consumer Discretionary 0 1 1 0 0
nike.com Consumer Discretionary 0 1 0 1 0
openai.com Information Technology 0 3 0 3 0
opentable.com Consumer Discretionary 0 2 0 2 0
oracle-mail.com Information Technology 0 5 1 4 0
oracle.com Information Technology 0 1 0 1 0
orbitz.com Consumer Discretionary 0 15 14 1 0
pandora.com Communication Services 0 36 17 19 0
pinterest.com Communication Services 0 33 1 32 0
planetfitness.com Consumer Discretionary 0 1 0 1 0
plex.tv Consumer Discretionary 0 4 0 4 0
pons.com Consumer Discretionary 0 1 0 1 0
pornhub.com Consumer Discretionary 0 3 0 3 0
quizlet.com Consumer Discretionary 0 5 0 5 0
quora.com Information Technology 0 66 0 63 3
realtor.com Consumer Discretionary 0 7 2 5 0
reddit.com Communication Services 0 1 0 1 0
redditmail.com Communication Services 4 85 0 65 20
reuters.com Communication Services 0 2 0 2 0
ring.com Consumer Discretionary 0 5 1 4 0
roblox.com Consumer Discretionary 0 1 0 1 0
roku.com Communication Services 0 3 0 3 0
scrimba.com Information Technology 0 17 12 4 1
shazam.com Communication Services 0 1 0 1 0
shein.com Consumer Discretionary 0 106 93 12 1
sheinnotice.com Consumer Discretionary 0 1 0 1 0
shopify.com Consumer Discretionary 0 1 0 1 0
shutterstock.com Consumer Discretionary 0 52 20 31 1
skillcrush.com Consumer Discretionary 0 27 12 15 0
snapchat.com Communication Services 0 5 0 5 0
soundcloud.com Communication Services 0 11 3 8 0
southwest.com Consumer Discretionary 0 1 0 1 0
springernature.com Consumer Discretionary 0 2 0 2 0
statista.com Information Technology 0 1 0 1 0
steampowered.com Consumer Discretionary 0 1 0 1 0
target.com Consumer Discretionary 4 109 87 22 0
temu.com Consumer Discretionary 0 3 0 3 0
temuemail.com Consumer Discretionary 0 61 40 20 1
thumbtack.com Consumer Discretionary 0 11 1 10 0
tiktok.com Communication Services 0 1 0 1 0
today.com Information Technology 4 116 112 3 1
topcoder.com Information Technology 0 1 0 1 0
tracfone.com Consumer Discretionary 0 10 5 5 0
trip.com Consumer Discretionary 0 2 0 2 0
tripadvisor.com Consumer Discretionary 0 74 32 42 0
trivago.com Consumer Discretionary 0 32 26 6 0
turo.com Consumer Discretionary 0 1 0 1 0
tutorialsdojo.com Information Technology 0 21 15 5 1
tutorialspoint.com Information Technology 0 23 16 7 0
twitch.tv Communication Services 0 6 0 6 0
uber.com Consumer Discretionary 4 173 169 4 0
udacity.com Information Technology 0 1 0 1 0
usnews.com Consumer Discretionary 0 2 0 2 0
vimeo.com Communication Services 0 14 10 4 0
walmart.com Consumer Discretionary 0 1 1 0 0
wayfair.com Consumer Discretionary 1 549 520 27 2
waze.com Information Technology 0 2 0 2 0
webmd.com Health 4 270 9 250 11
wish.com Consumer Discretionary 1 501 455 46 0
xhamster.com Consumer Discretionary 0 1 0 1 0
xvideos.com Consumer Discretionary 0 1 0 0 1
yelp.com Consumer Discretionary 0 11 0 11 0
zillow.com Consumer Discretionary 0 1 0 1 0